OpenStudy (anonymous):

On problem set 1, can anyone explain supplemental problem 1A6 and how to approach it?

OpenStudy (phi):

Do you mean the question that starts, "A small plane wishes to fly due north at 200 mph..."?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

create a [N E] or [x y] vector to represent the plane's velocity and do the same for the wind.

OpenStudy (phi):

For example, let the vector have components be i and j, as asked for in the question +i will mean east, +j will mean north (negative numbers will mean west or south) using that definition , the velocity vector will have length 200 and direction north. in other words, write plane= [ 0 200 ] This will be the speed measured over ground

OpenStudy (phi):

the wind is *from* the northeast, so it blows towards the southwest this direction can be represented by a vector pointing to the left 1 unit and down 1 unit [-1 -1] make it "unit length" by dividing by its magnitude= \( \sqrt{(-1)^2 + (-1)^2}=\sqrt{2} \) [-1 -1]/sqr(2) now "scale" by 50 to make this vector the correct length wind= (50/sqr(2) ) * [-1 -1]

OpenStudy (phi):

Here is a graph showing what we want